hsdfhsfd:

Which is the correct way to model the equation Negative x + 8 = 7 x + (negative 8) using algebra tiles? 1 positive x-tile and 8 positive unit tiles on the left side; 7 positive x-tiles and 8 positive unit tiles on the right side 8 positive unit tiles on the left side; 7 positive x-tiles and 8 negative unit tiles on the right side 1 negative x-tile and 8 positive unit tiles on the left side; 7 positive x-tiles and 8 negative unit tiles on the right side 8 positive x-tiles and 1 negative unit tile on the left side; 8 negative x-tiles and 7 positive unit tiles on the right side

hsdfhsfd:

plz help

Hoodmemes:

In order to have both sides equal to each we must have terms that be canceling out: First step: \[x + 8 = 7\times x + 8\]

Hoodmemes:

Second step: \[8 \times x= 7\times x-8\]

Hoodmemes:

Third step: \[-x + 8=7 \times x -8\]

Hoodmemes:

Forth step: \[8\times x -1=7-8\times x\] And the answer would be in the (blank) step because it gives an answer of x=8 when solved